The world is risk neutral and interest rates are 20%. With probability ¼, your firm will be worth $60 next year. With probability ¾, it will be worth $100. What interest rate do you have to promise to raise $70 in debt today?
Answer #1

Let the rate to be promised be r

70*(1+20%)=1/4*60+3/4*(70*(1+r))
=>r=92/70-1
=>r=31.429%